OpenStudy (anonymous):

Find the common difference for the sequence shown. 56, 49, 42, 35, ... -12 -7 7

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

What is the method for finding common difference?

OpenStudy (anonymous):

The difference is 7. You cant have a difference of a negative number.

OpenStudy (jhannybean):

56 - 49 = ? 49 - 42 = ? 42 - 35 = ? The difference is found when you subtract the first number from the one coming after it.

OpenStudy (anonymous):

yes the difference is 7 but you have to use a negative to subtract otherwise you would be adding.

OpenStudy (anonymous):

Common difference is not first number - second number. It is second number - first number. It can be positive or negative, which denotes whether the sequence is increasing or decreasing. The common difference is 49 - 56 = -7

OpenStudy (jhannybean):

Ahh, my bad.

OpenStudy (ybarrap):

$$\tt a_1+(n-1)d=a_n\text{ for n=1,2,... and d>0. }\\\text{Here d = 7}. $$

OpenStudy (anonymous):

@ybarrap If you put 7 for d in your equation you will get a sequence different from the one in the question. d can be negative as well.

OpenStudy (ybarrap):

You are right! d = -7, sorry about that. $$a_1=56\\ a_2=56+(2-1)*(-7)=56-7=49\\ a_3=56+(3-1)*(-7)=56-14=42\\ etc. $$